OpenStudy (anonymous):

A 40 foot beam is to be cut in two. If one portion is one third the length of the other portion, what are the two lengths of the portions?

OpenStudy (ghazi):

let's say length of one portion is x then remaining portion would be 40-x . now according to the question x=1/3*(40-x) now solve for x

OpenStudy (campbell_st):

the beam is to be cut in the ratio 1:3.. use the unitary method to find the lengths find the length of 1 part in the ratio How many parts in the ratio: 1 + 3 divide the length by the number of parts in the ratio... this will give the shorter length... triple it to get the longer length. To check you answer... add the 2 lengths and you should have 40 feet

OpenStudy (anonymous):

x=52

OpenStudy (campbell_st):

nope.... sorry

OpenStudy (campbell_st):

isn't 52 longer than your beam

OpenStudy (anonymous):

i dont understand what i'm doing

OpenStudy (campbell_st):

ok... look at my explanation cutting it in a ratio 1:3 what is the sum of the parts of the ratio...?

OpenStudy (anonymous):

x/3 + x =40 x+3x=120 4x=120 X=30 x/3=10

OpenStudy (campbell_st):

correct

OpenStudy (campbell_st):

a simpler way is ratio 1:3 to 4 parts 40/4 = 10 the shorter part is 10 feet the longer is 3 x 10 = 30 feet this is called the unitary method..